# Sparse Autoencoders Can Interpret Randomly Initialized Transformers

### v1

@@ 4 A toy model of superposition
Figure 4: An illustrative example of the effect of a randomly initialized neural network on superposed input data. We take 10K samples of ${n_{s}}=3$ sparse input features from a Lomax distribution with shape $\alpha=1$ and scale $\lambda=1$ and project these to ${n_{d}}=2$ dense input features by an IID standard normal matrix. Then, we pass the dense outputs to a two-layer MLP with ReLU activation and hidden size of $4{n_{d}}$ and recover ${n_{s}}=3$ sparse outputs by the inverse of the previously generated projection matrix.
@@ 4.1 Matrix multiplications preserve superposition
First, we consider a simplified model to demonstrate that multiplication by a weight matrix $W$ preserves superposition. Imagine that we generate superposed input data $x$ by first generating ${n_{s}}$ IID ‘sparse’ features $z$ from a heavy-tailed Lomax distribution $z\sim\operatorname{Lomax}(\alpha,\lambda)$. We can project the higher-dimensional, sparse $z$ down to lower-dimensional, dense $x$ with a matrix $D$, then add Gaussian noise with a small variance $\Sigma$,
$x\sim\mathcal{N}(x;Dz,\Sigma).$ (1)
Importantly, if we multiply $x$ by some matrix $W$, then $x^{\prime}=Wx$ is also superposed: it is generated by the same model as $x$, except with different noise covariances and mappings from $z$ to $x$, namely
$x^{\prime}\sim\mathcal{N}(z;WDz,W\Sigma W^{T}).$ (2)
We can see that the same intuition might extend to neural networks with non-linearities by visualizing the results of passing the dense activations through a simple feed-forward network (MLP). Figure 4 shows an example where ${n_{s}}=3$ sparse features are projected down to ${n_{d}}=2$ dense features, where the MLP outputs appear superposed despite the non-linearity. Moreover, it suggests that NNs might amplify superposition rather than only preserving it: comparing the inputs (Figures 4(a) and 4(b)) to the outputs (Figures 4(d) and 4(c)), there are fewer points between the ‘arms’ of the outputs.
@@ 4.2 Do random NNs preserve or amplify superposition?
We investigated more systematically by generating toy data by the same procedure as Sharkey et al. (2022), i.e., sampling ground-truth features on a hypersphere and generating correlated feature coefficients such that only a small number are active (Appendix C.1). We then passed these inputs to a two-layer MLP at initialization and trained SAEs on both the inputs and outputs individually. As a control, we used Gaussian-distributed inputs with a mean and standard deviation equal to the superposed toy data. We used standard SAEs with an $L^{1}$ sparsity penalty (Appendix C.2).
@@
Following Sharkey et al. (2022), we confirmed that SAEs can recover the ground-truth features that generated the data (Figure 5). In particular, we measured the mean max cosine similarity (MMCS): for every data-generating feature, we found its maximum cosine similarity with the features learned by the SAE (its decoder weight vectors) and took the average over data-generating features. However, the MMCS only applies to the MLP inputs, where we have access to the data-generating features – a different approach is required to analyze the MLP outputs. To this end, we took the ability of SAEs to achieve low reconstruction error with high sparsity as a proxy for the degree to which the training data exhibits superposition. Specifically, we vary the $L^{1}$ penalty coefficient to obtain Pareto frontiers of the explained variance against sparsity measures (Figure 6).
As expected, we found that SAEs achieved much greater sparsity at a given level of explained variance for the superposed inputs relative to the Gaussian control (Figure 6, light blue and green). Interestingly, the difference between the superposed outputs, i.e., the outputs of the MLP given the superposed inputs, and the Gaussian outputs is much smaller, with only slightly greater sparsity at a given level of explained variance. This suggests that the outputs of randomly initialized MLPs have a relatively high level of sparsity insensitive to the input distribution. We consider other sparsity measures and hyperparameters in Appendix C.
Figure 6: Pareto frontiers of the explained variance (normalized reconstruction error) against the $L^{1}$ norm divided by the square root of the $L^{2}$ norm (sparsity) for toy datasets generated to exhibit superposition, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP (Section 4.2). Each point denotes a choice of $L^{1}$ penalty coefficient, and the error bars denote the standard error in the mean taken over random seeds.
@@ 4.3 Do token embeddings exhibit superposition?
Figure 7: Pareto frontiers of the explained variance (normalized reconstruction error) against the $L^{1}$ norm divided by the square root of the $L^{2}$ norm (sparsity) for $300$-dimensional GloVe word vectors (Pennington et al., 2014), Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP (Section 4.2). Each point denotes a choice of $L^{1}$ penalty coefficient.
We find that there is a smaller gap between the Pareto frontiers of the GloVe word vectors and the corresponding Gaussian controls (Figure 7) than in the case of the toy superposed datasets described in Section 4.2 (Figure 6). More interestingly, we again see that the Pareto frontiers for both inputs improve when they are passed to a randomly initialized two-layer MLP, emphasizing the possibility that random NNs ‘sparsify’ their inputs (i.e., increase the degree to which superposition is apparently exhibited).
@@ 5 Conclusion
We have shown that when SAEs are applied to transformer language models whose parameters are randomized, the SAE latents are approximately equally interpretable to when SAEs are applied to trained language models, measured in terms of auto-interpretability.
This result shows that while SAEs learn apparently interpretable latents, this does not necessarily imply that SAEs discover the computational structure of the underlying model. Instead, it may be the case that the interpretable latents are a result of the inherent sparsity of the text data on which language models are trained, which is preserved or even amplified by neural networks, regardless of whether the network is trained or randomized. Additionally, it may be that current evaluations fail to capture the “abstractness” of different SAE latents.
Our findings raise important questions about how to understand results based on SAEs; in particular, they highlight the need to develop better techniques to analyze the computational structure learned by language models. We would also recommend that future work in mechanistic interpretability benchmarks its techniques against randomly initialized models as a control to properly attribute structure to what is learned when training the underlying model.
Appendix A Auto-interpretability ROC curves
Figures  8,  9,  10 show the similarity between AUROC for the trained and randomized SAEs for the 70M, 160M and 1B models.
@@ Appendix A Auto-interpretability ROC curves
Figure 10: ROC curves for ‘fuzzing’ auto-interpretability for Pythia-1B over 100 SAE latents. These results demonstrate the similarity in performance between the SAE variants, although here we do not observe an overall degredation in quality.
Appendix B Hyperparameter sweep for 160M models
Here we demonstrate the stability of our results over a hyperparamter sweep of the expansion factor and $k$ for Top-k in the SAEs trained on the Pythia-160M model.
@@ Appendix B Hyperparameter sweep for 160M models
Figure 11: Robustness of sparse autoencoder performance to hyperparameter selection. Performance metrics remain stable across a wide range of expansion factors (16-128) and sparsity parameters k (16-32), with all initialization strategies maintaining their relative performance ordering. This stability suggests that moderate hyperparameter values (e.g., expansion factor=64, k=32) are sufficient.
Appendix C A toy model of superposition
@@
In Section 4, we trained SAEs on toy data designed to exhibit superposition (Sharkey et al., 2022) and GloVe word vectors (Pennington et al., 2014). In this section, we detail the data-generation procedure and training setup.
C.1 Data generation
@@
Like Sharkey et al. (2022), we choose ${n_{s}}=512$, ${n_{d}}=256$, $\lambda=0.99$, and $m=5$. We include a Python implementation of this procedure in Figure 12.
C.2 Training
@@ 9 Args:
Figure 12: A Python implementation of the data-generation procedure introduced by Sharkey et al. (2022) and used in Section 4.
Figure 13: Pareto frontiers of the explained variance (normalized reconstruction error) against the $L^{0}$ norm (sparsity) for toy datasets generated to exhibit superposition, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Figure 14: Pareto frontiers of the explained variance (normalized reconstruction error) against the $L^{1}$ norm (sparsity) for toy datasets generated to exhibit superposition, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Figure 15: Pareto frontiers of explained variance (normalized reconstruction error) against the $L^{1}$ norm over the square root of the $L^{2}$ norm (sparsity) for toy datasets generated to exhibit superposition, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Figure 16: Pareto frontiers of explained variance (normalized reconstruction error) against the Hoyer sparseness (sparsity) for toy datasets generated to exhibit superposition, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Figure 17: Pareto frontiers of explained variance (normalized reconstruction error) against sparsity measures for 50-dimensional GloVe word vectors, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Figure 18: Pareto frontiers of explained variance (normalized reconstruction error) against sparsity measures for 100-dimensional GloVe word embeddings, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Figure 19: Pareto frontiers of explained variance (normalized reconstruction error) against sparsity measures for 200-dimensional GloVe word embeddings, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Figure 20: Pareto frontiers of explained variance (normalized reconstruction error) against sparsity measures for 300-dimensional GloVe word embeddings, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Appendix D Example features for Pythia-6.9B
Here we provide some example features for the SAEs trained on the Pythia-6.9B model.
@@
    Text: ¡—endoftext—¿. If for any reason you are unhappy with our service please contact us directly so we can make it right for you.Journal of Cyber Security, Vol.
@@
    Text: ¡—endoftext—¿ Security Practitioner. Tremendously passing CompTIA Advanced Security Practitioner (casp) cert has never been as easy as it
@@
[…]
@@
Variant: Randomized excl emb
@@
Variant: Randomized incl emb
@@
    Text: ¡—endoftext—¿ to the 21st Century. Rhodes won the Pulitzer prize for The Making of the Atomic Bomb (01987) his first of four books chronicling the
@@
    Text: ¡—endoftext—¿, James Coburn, movies we love, Pulp Consumption, Steve McQueen, Western, Yul Brynner. Bookmark the permal
@@
    Text: ¡—endoftext—¿ance - Chapters: 1 - Words:. Fruits Basket - Rated: T - English - Romance/Angst - Chapters:

### v2

@@ 4 A toy model of superposition in random networks
Figure 3: An example of the effect of a randomly initialized neural network on superposed input data. We take 10K samples of ${n_{s}}=3$ sparse input features from a Lomax distribution with shape $\alpha=1$ and scale $\lambda=1$ and project these to ${n_{d}}=2$ dense input features by an i.i.d. standard normal matrix. Then, we pass the dense outputs to a two-layer MLP with ReLU activation and hidden size of $4{n_{d}}$ and recover ${n_{s}}=3$ sparse outputs by the inverse of the previously generated projection matrix.
@@ 4.1 Matrix multiplications preserve superposition
First, we consider a simplified model to demonstrate that multiplication by a weight matrix $W$ preserves superposition. Imagine that we generate superposed input data $x$ by first generating $n_{s}$} i.i.d. ‘sparse’ features $z$ from a heavy-tailed Lomax distribution $z\sim\operatorname{Lomax}(\alpha,\lambda)$. We can project the higher-dimensional, sparse $z$ down to lower-dimensional, dense $x$ with a matrix $D$, then add Gaussian noise with a small variance $\Sigma$, $x\sim\mathcal{N}(x;Dz,\Sigma)$. Importantly, if we multiply $x$ by some matrix $W$, then $x^{\prime}=Wx$ is also superposed: it is generated by the same model as $x$, except with different noise covariances and mappings from $z$ to $x$, namely $x^{\prime}\sim\mathcal{N}(z;WDz,W\Sigma W^{T})$.
We can see that the same intuition might extend to neural networks with nonlinearities by visualizing the results of passing the dense activations through a simple feed-forward network (MLP). Figure 3 shows an example where ${n_{s}}=3$ sparse features are projected down to ${n_{d}}=2$ dense features, and the MLP outputs appear superposed despite the non-linearity. Moreover, it suggests that NNs might amplify superposition rather than only preserving it: comparing the inputs (Figures 3(a) and 3(b)) to the outputs (Figures 3(d) and 3(c)), there are fewer points between the ‘arms’ of the outputs.
@@ 4.2 Do random NNs preserve or amplify superposition?
We investigated this suggestion by generating toy data with the same procedure as Sharkey et al. (2022), i.e., sampling ground-truth features on a hypersphere and generating correlated feature coefficients such that only a small number are active (Appendix I.1). We then passed these inputs to a two-layer MLP at initialization and trained SAEs on both the inputs and outputs individually. As a control, we used Gaussian-distributed inputs with a mean and standard deviation equal to the superposed toy data. We used standard SAEs with an $L^{1}$ sparsity penalty (Appendix I.2).
@@
Following Sharkey et al. (2022), we confirmed that SAEs can recover the ground-truth features that generated the data (Figure 4). In particular, we measured the mean max cosine similarity (MMCS): for every data-generating feature, we found its maximum cosine similarity with the features learned by the SAE (its decoder weight vectors) and took the average over data-generating features. However, the MMCS only applies to the MLP inputs, where we have access to the data-generating features – a different approach is required to analyze the MLP outputs. To this end, we took the ability of SAEs to achieve low reconstruction error with high sparsity as a proxy for the degree to which the training data exhibits superposition. Specifically, we vary the $L^{1}$ penalty coefficient to obtain Pareto frontiers of the explained variance against sparsity measures (Figure 5(a)).
As expected, we found that SAEs achieved much greater sparsity at a given level of explained variance for the superposed inputs relative to the Gaussian control (Figure 5(a); orange and blue-green). Interestingly, the difference between the superposed outputs, i.e., the outputs of the MLP given the superposed inputs, and the Gaussian outputs is much smaller, with only slightly greater sparsity at a given level of explained variance. This suggests that the outputs of randomly initialized MLPs have a relatively high level of sparsity insensitive to the input distribution. We consider other sparsity measures and hyperparameters in Appendix I.
(a) Toy datasets following Sharkey et al. (2022).
(b) 300300-dim. GloVe vectors (Pennington et al., 2014).
Figure 5: Pareto frontiers of the explained variance against the $L^{1}$ norm divided by the square root of the $L^{2}$ norm (sparsity) for datasets perhaps exhibiting superposition, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP (Section 4.2. Each point denotes a choice of $L^{1}$ penalty coefficient.
@@ 4.3 Do token embeddings exhibit superposition?
We find that the gap between the Pareto frontiers of the GloVe word vectors and the corresponding Gaussian controls (Figure 5(b)) is smaller than that observed for the toy superposed datasets described in Section 4.2 (Figure 5(a)). More interestingly, we again see that the Pareto frontiers for both inputs improve when they are passed to a randomly initialized two-layer MLP, emphasizing the possibility that random NNs ‘sparsify’ their inputs (i.e., increase the degree of apparent superposition).
5 Limitations
In this work, we demonstrate that auto-interpretability measures can produce apparently meaningful, interpretable results for SAEs trained on randomly initialized models, which are unlikely to exhibit computationally interesting features. Given the impossibility of testing across all datasets and model architectures, we strategically focused on the Pythia family of models, widely adopted in mechanistic interpretability research (e.g., Paulo and Belrose, 2025; Ghilardi et al., 2025; Mueller, 2024), and the RedPajama-V2 dataset, representing typical pre-training data for language models and SAEs.
While we used the default model for generating explanations in the EleutherAI auto-interpretability framework (Caden et al., 2025), exploring alternative models could yield valuable insights into aggregate behaviors and the quality of generated explanations. Importantly, we do not claim that SAEs fail to capture information from trained Transformers above and beyond randomly initialized transformers; only that aggregate auto-interpretability measures do not necessarily indicate the existence of interesting underlying features.
@@ 6 Conclusion
In this work, we applied sparse autoencoders to both trained and randomly initialized transformers and evaluated them with a suite of common quantitative metrics. Our central empirical finding is that, under certain conditions, these metrics – particularly aggregate auto-interpretability scores – can be surprisingly similar in both settings. While we observe that features derived from trained transformers are qualitatively more complex and abstract, especially in later layers, these aggregate metrics often fail to capture this distinction.
This result does not imply that SAEs trained on real models fail to learn meaningful computational features. Rather, it reveals a limitation in our current evaluation methods. High aggregate auto-interpretability scores are insufficient proof for the discovery of complex, learned computations: they may instead reflect simpler structure inherent in the data or model architecture that is preserved even by random weights. Our analysis of token distribution entropy, while preliminary, serves as a proof-of-concept: it successfully revealed differences in feature ‘abstractness’ that aggregate auto-interpretability scores missed. Future work should focus on developing more robust metrics that can quantify the computational significance of the features SAEs discover. Our work reaffirms the importance of benchmarking interpretability techniques against strong, appropriately constructed null models, such as the randomly initialized transformers used here. Without such baselines, it is difficult to confidently attribute discovered features to the process of learning.
Appendix A Broader Impact
This work investigates a method currently used for mechanistic interpretability of LLMs, yielding results that challenge certain assumptions about sparse autoencoders. By demonstrating that SAEs can produce similar aggregate auto-interpretability scores for both random and trained transformers, our findings raise important questions about what these SAE evaluation methods are actually capturing.
By better understanding the metrics of SAE quality, we hope that this work will contribute to a more informed search of better SAE-like methods and thus help to make these models more interpretable and to mitigate the potential harm these models could cause. Since our work is an empirical study of the capabilities of a presently used method, and it shows that the method provides interpretation of both random and trained transformers, we think the risk that this work could lead to negative social impact is minimal.
Appendix B Auto-interpretability ROC curves
Figures  6,  8,  12 show the similarity between ‘fuzzing’ AUROC for the trained and randomized SAEs for the 70M, 160M, and 1B models. Figures  7,  9,  13, show the similarity between ‘detection’ AUROC for the trained and randomized SAEs for the 70M, 160M, and 1B models.
B.1 Pythia 70m
@@
Figure 7: ROC curves for ‘detection’ auto-interpretability for Pythia-70m over 100 SAE latents. These results demonstrate the similarity in performance between the SAE variants, as well as the overall degradation in performance as the layer index increases.
B.2 Pythia 160m
@@
Figure 9: ROC curves for ‘detection’ auto-interpretability for Pythia-160m over 100 SAE latents. These results demonstrate the similarity in performance between the SAE variants, as well as the overall degradation in performance as the layer index increases.
B.3 Pythia 410m
Figure 11: ROC curves for ‘detection’ auto-interpretability for Pythia-410m over 100 SAE latents. These results demonstrate the similarity in performance between the SAE variants, as well as the overall degradation in performance as the layer index increases.
B.4 Pythia-1b
Figure 12: ROC curves for ‘fuzzing’ auto-interpretability for Pythia-1b over 100 SAE latents. These results demonstrate the similarity in performance between the SAE variants, although here we do not observe an overall degradation in quality.
Figure 13: ROC curves for ‘detection’ auto-interpretability for Pythia-1b over 100 SAE latents. These results demonstrate the similarity in performance between the SAE variants, although here we do not observe an overall degradation in quality.
B.5 Pythia 6.9b
Figure 14: ROC curves for ‘detection’ auto-interpretability for Pythia-6.9b over 100 SAE latents. These results demonstrate the similarity in performance between the SAE variants.
Appendix C Effect of increased training data
For our primary experiments, we trained SAEs on 100M tokens (Section 3). We verified that our results were not explained by a lack of sufficient training data by repeating a subset of these experiments with SAEs trained on 1B tokens from the RedPajama dataset (Figure 15).
Figure 15: Evaluation metrics for SAEs trained with one billion tokens on the Pythia-70m and 410m models. These results correspond to columns of Figure 2, which show the same evaluation metrics for SAEs trained on 100M tokens, and qualitatively similar behavior.
Appendix D Effect of decreased training data for Pythia-1B
Figure 16: Evaluation metrics for SAEs trained with 1M and 1B tokens on Pythia-1b. The explained variance and CE loss score are significantly lower for the 1M model, showing that the SAEs are under-trained. Average auto-interpretability scores are slightly lower for the earliest layers, but decline sharply with increasing layer. The trends in auto-interpretability and token distribution entropy with layer index are consistent with other SAEs.
Appendix E Uncertainty plots for Pythia-70m
We computed uncertainty for our evaluation metrics on Pythia-70m using five random seeds.
Figure 17: Uncertainty for Pythia-70m metrics computed using five random seeds.
Appendix F Effect of SAE hyperparameters for Pythia-160m
@@
Figure 18: Robustness of SAE performance to hyperparameter selection. Standard evaluation metrics remain stable across a wide range of expansion factors $R$ (16 to 128) and sparsities $k$ (16 to 32), with all initialization strategies maintaining their relative performance ordering. This stability suggests that moderate hyperparameter values (e.g., expansion factor $R=64$, sparsity $k=32$) suffice.
Appendix G Effect of SAE hyperparameters for Pythia-1b
Figure 19: Evaluation metrics for SAEs trained on the Pythia-1b model with different hyperparameters, including the main results from Figure 2. SAEs with a very small expansion factor $R=2$ and sparsity $k=4$ are clearly distinguished from our default hyperparameters by the explained variance and CE loss score. Importantly, the auto-interpretability scores of these SAEs remain similar to those trained with default hyperparameters on either trained or randomised models.
Appendix H Token distribution entropy vs. auto-interpretability
Figure 20: Scatter plots of the per-latent token distribution entropy against ‘fuzzing’ AUROC (auto-interpretability score) for SAEs trained on multiple layers of the Pythia-6.9b model. Each point corresponds to a single latent, taken from the sample of latents used to compute the aggregate metrics displayed in Figures 1 and 2.
Figure 20 clearly distinguishes the negative control, randomized variants, and the trained variant:
  • Control: Latents have a consistently high entropy (i.e., max activating examples with activation patterns spread across many tokens) and low auto-interpretability score (i.e., generated explanations that fail to adequately explain these activation patterns). No correlation between the two variables is evident.
  • Randomized: For each of the randomization schemes described in Section 3, we see a negative correlation between entropy and auto-interpretability: in general, the wider variety of tokens for which a latent is activated, the less well the latent’s activation patterns are explained by its generated explanation.
  • Trained: There is a weaker correlation between the two variables. Crucially, in addition to the broad trend observed for the randomized variants, we also see latents with high entropy and auto-interpretability. Some latents have activation patterns that are spread across multiple tokens, which are nevertheless consistent with the latent’s generated explanation.
These results are consistent with the view that aggregate auto-interpretability scores obscure the differences between SAEs based on trained and randomized models. While randomized models with consistent token embeddings can produce ‘single-token’ features, whose activation patterns are easy to explain, only Transformers trained on natural language produce more complex semantic features.
Appendix I A toy model of superposition
@@
In Section 4, we trained SAEs on toy data designed to exhibit superposition [Sharkey et al., 2022] and GloVe word vectors [Pennington et al., 2014]. In this section, we detail the data-generation procedure and training setup.
I.1 Data generation
@@
Like Sharkey et al. [2022], we choose ${n_{s}}=512$, ${n_{d}}=256$, $\lambda=0.99$, and $m=5$. We include a Python implementation of this procedure in Figure 21.
I.2 Training
@@ 9 Args:
Figure 21: A Python implementation of the data-generation procedure introduced by Sharkey et al. [2022] and used in Section 4.
Figure 22: Pareto frontiers of the explained variance against the $L^{0}$ norm (sparsity) for toy datasets generated to exhibit superposition, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Figure 23: Pareto frontiers of the explained variance against the $L^{1}$ norm (sparsity) for toy datasets generated to exhibit superposition, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Figure 24: Pareto frontiers of explained variance against the $L^{1}$ norm over the square root of the $L^{2}$ norm (sparsity) for toy datasets generated to exhibit superposition, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Figure 25: Pareto frontiers of explained variance against the Hoyer sparseness (sparsity) for toy datasets generated to exhibit superposition, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Figure 26: Pareto frontiers of explained variance against sparsity measures for 50-dimensional GloVe word vectors, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Figure 27: Pareto frontiers of explained variance against sparsity measures for 100-dimensional GloVe word embeddings, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Figure 28: Pareto frontiers of explained variance against sparsity measures for 200-dimensional GloVe word embeddings, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Figure 29: Pareto frontiers of explained variance against sparsity measures for 300-dimensional GloVe word embeddings, Gaussian controls with the same mean and variance, and the corresponding outputs when these are passed to a randomly initialized two-layer MLP.
Appendix J Example features for Pythia-6.9b
@@
    Text: <endoftext—¿—. If for any reason you are unhappy with our service please contact us directly so we can make it right for you.Journal of Cyber Security, Vol.
@@
    Text: <endoftext—¿— Security Practitioner. Tremendously passing CompTIA Advanced Security Practitioner (casp) cert has never been as easy as it
@@
    Text: Here, we introduce the coordinate systems for three-dimensional space $\square\square\square$2. The study of 3-dimensional spaces lead us to the setting for our study
@@
Variant: Randomized excluding embeddings
@@
Variant: Randomized including embeddings
@@
    Text: <endoftext—¿— to the 21st Century. Rhodes won the Pulitzer prize for The Making of the Atomic Bomb (01987) his first of four books chronicling the
@@
    Text: <endoftext—¿—, James Coburn, movies we love, Pulp Consumption, Steve McQueen, Western, Yul Brynner. Bookmark the permal
@@
    Text: <endoftext—¿—ance - Chapters: 1 - Words:. Fruits Basket - Rated: T - English - Romance/Angst - Chapters:
@@
Appendix K Compute Details
We performed all experiments with a single NVIDIA A100 80GB GPU in a private cluster. Table 1 lists the approximate duration of the final experiments for each model size and transformer variant. We estimate that the total cost of preliminary and failed experiments is roughly equal to the cost of the final experiments.
Model Variants Approx. time per variant (hours) Total time (hours)
Pythia-6.9b 5 70 350
Pythia-1b 5 10 50
Pythia-410m 5 5 25
Pythia-160m 5 1 5
Pythia-70m 5 1 5
Overall time: 435
Table 1: Approximate time required for our experiments.
Appendix L Example feature dashboards for Pythia-6.9b
Here we provide more detailed ‘feature dashboards,’ including per-feature activation patterns, token distribution entropy, and auto-interpretability (‘fuzz’ ROC) scores. We include two randomly sampled features for the control, randomized, and trained variants described in Section 3, trained on every fourth layer of Pythia-6.9b.
L.1 Trained
L.2 Control
L.3 Randomized excluding embeddings
L.4 Randomized including embeddings
L.5 Step 0